# Addressing Data Heterogeneity in Federated Learning with Adaptive Normalization-Free Feature Recalibration

### v1

@@ 5 Conclusion
We introduce ANFR, a simple approach combining the strengths of weight standardization and channel attention to address the challenges of data heterogeneity and privacy at a design level in FL. ANFR fills a gap by being the first method to simultaneously work in GFL, pFL, and private FL scenarios while being compatible with any aggregation method and offering a robust increase in performance. Extensive experiments demonstrate the superior adaptability and performance of ANFR, as it consistently surpasses the performance of baseline architectures, regardless of the aggregation method employed. Our results position ANFR as a compelling backbone model suitable for both global and personalized FL scenarios where statistical heterogeneity and privacy guarantees are important concerns. Our findings highlight the need to look beyond aggregation methods as the core component of federated performance and the critical role of architectural innovations in reaching the next frontier in private and collaborative settings.
Reproducibility Statement
We have taken several steps to ensure reproducibility of our results. Details for the datasets we used and their pre-processing are in Appendix A.1. We also introduce a dataset of our own creation, FedChest, and provide detailed instructions on how to reproduce it in Appendix B. How we tuned our models is explained in Appendix A.2. This paper introduces a model, so we will provide ImageNet pre-trained weights for it upon acceptance. Details for all the models we use are in Appendix A.3. We will open-source all of our code upon acceptance.
@@ A.1 Datasets
Chest X-Ray Multi-Label Classification on FedChest. Please refer to Appendix B.1.
@@ A.2 Hyper-parameter Tuning
A.3 Model Details
Table 6 presents pre-training details, parameter counts, multiply-accumulate counts (GMACs) and floating point operation counts (FLOPs) and ImageNet (Russakovsky et al., 2015) validation set top-1 performance for all of the compared models. For models which are pre-trained by us, links to the pre-trained weights will be made public after acceptance.
We note that the computational overhead of adding channel attention modules—which is already small—is offset by the abolishment of BN which is more expensive than SWS and incurs a bigger memory and time overhead (Brock et al., 2021b).
| Model | Parameters | GMACs | GFLOPs | IN-1K performance | Training Recipe |
@@ A.3 Model Details
Appendix B FedChest construction and additional results
B.1 Construction and hyper-parameters
@@
After pre-processing, CheXpert has twice as many samples as the other datasets, so we further split it into two clients, cxp_young and cxp_old using the median age of the patient population (63 years), leading to a total of 4 clients with train/val/test splits of (given in thousands): for CXR14, for PadChest, for cxp_old and . The task is multi-label classification across the 8 common classes.
@@ B.2 Additional Results
Appendix C Tabular Comparison with Related Work

### v2

@@ 5 Conclusion
We introduce ANFR, a novel design-level approach to address the challenges of data heterogeneity at a design level in FL. ANFR fills a research gap by simultaneously working in GFL, pFL, and private FL scenarios while being compatible with any aggregation method and offering a robust increase in performance. Extensive experiments demonstrate the superior adaptability and performance of ANFR, as it consistently surpasses the performance of baseline architectures, regardless of the aggregation method employed. Our results position ANFR as a compelling backbone model suitable for both global and personalized FL scenarios where statistical heterogeneity and privacy guarantees are important concerns. Our findings highlight the need to look beyond aggregation methods as the core component of federated performance and the critical role of architectural innovations in reaching the next frontier in private and collaborative settings.
@@ A.1 Datasets
Chest X-Ray Multi-Label Classification on FedChest. Please refer to Appendix C.1.
@@ A.2 Hyper-parameter Tuning
A.3 Model Details and Computational Overhead
Table 10 presents pre-training details, parameter counts, multiply-accumulate counts (GMACs) and floating point operation counts (FLOPs) and ImageNet (Russakovsky et al., 2015) validation set top-1 performance for all models. For models which are pre-trained by us, links to the pre-trained weights will be made public after acceptance.
Model | Parameters | GFLOPs | GMACs | IN-1K performance | Training Recipe |
@@ A.3 Model Details and Computational Overhead
Appendix B Additional Results
B.1 Results on Fed-ISIC2019 using FLamby hyper-parameters
The experimental setup we use for Fed-ISIC2019 in the main paper is an improved version of the example benchmark presented in Section 4.1 of Ogier du Terrail et al. (2022), so one might wonder how the compared models perform under the original settings. To answer this we repeat Centralized, FedAvg, and SCAFFOLD training on Fed-ISIC2019 after aligning our hyper-parameters with Ogier du Terrail et al. (2022), meaning we reduce local steps to 100 without a scheduler, perform 9 federated rounds, and use pre-computed class weights in the focal loss. Results are presented in Table 11, showing ANFR comprehensively beats competing baselines, with an even wider performance gap compared to our original setting. The overall level of performance, including the gap between centralized and FL training, aligns with the results presented in Ogier du Terrail et al. (2022), as we expect. Additionally, SE-ResNet performs better than ANFR in centralized training, but the opposite is true in FL training, further validating our core claims in Section 3 that CA needs Weight Standardization to optimally adjust channel responses in heterogeneous FL. Although these results further support our claims, we believe the optimized version of Fed-ISIC2019 training we provide in the main paper is more of use to the community.
BN-ResNet | GN-ResNet | SE-ResNet | NF-ResNet | ANFR (Ours) | |
FedAvg | 59.50.75 | 55.262.96 | 61.921.58 | 60.760.75 | 65.341.29 |
SCAFFOLD | 57.612.78 | 57.622.95 | 67.340.42 | 57.350.73 | 71.071.27 |
Central | 61.262.92 | 57.091.85 | 73.001.09 | 61.281.53 | 72.031.55 |
B.2 Results Using Randomly Initialized Models
Given the ubiquity and demonstrated utility of ImageNet pre-trained models in FL (Qu et al., 2022; Pieri et al., 2023; Siomos et al., 2024), we use pre-trained models in the main paper. Nevertheless, we conduct additional experiments with FedAvg on CIFAR-10, FedChest and Fed-ISIC2019, using randomly initialized models. Although the results below bolster our claims, we avoided this setting initially as random weight initialization is not representative of the current standard settings adopted by FL practitioners. The only changes made to accommodate the absence of pre-training are to change the optimizer to AdamW and the learning rate to for CIFAR-10, and to double the number of local steps for Fed-ISIC2019. Our results in Table 12 show the same trend, of a gap existing between FL and centralized training but being smaller when using pre-trained models. In this setting, too, ANFR is the best performer.
Dataset | CIFAR-10 | Fed-ISIC2019 | FedChest | ||
Model | FedAvg | Central | FedAvg | FedAvg | Central |
BN-ResNet | 80.89 | 89.05 | 54.02 | 78.44 | 82.58 |
GN-ResNet | 78.52 | 86.69 | 54.92 | 73.68 | 80.82 |
SE-ResNet | 81.19 | 88.65 | 53.20 | 78.79 | 82.16 |
NF-ResNet | 81.66 | 88.96 | 56.75 | 79.06 | 83.55 |
ANFR (Ours) | 83.20 | 89.58 | 57.71 | 79.41 | 83.67 |
B.3 CIFAR-10 experiment without early-stopping
The results presented in Section 4.2 follow the experimental set-up of Pieri et al. (2023), where the validation set is used a form of early stopping in the following way: at every round the performance on the test set is only evaluated if the accuracy on the validation set has increased. While this is a methodologically valid set-up, it is also interesting to see how the models perform when no early-stopping is used. To compensate for this and avoid overfitting we disable gradient clipping and increase the batch size to 64. The results are presented in Table 13, showing how ANFR continues to beat the baselines.
Model | BN-ResNet | SE-ResNet | GN-ResNet | NF-ResNet | ANFR |
FedAvg | 67.39 | 74.75 | 96.73 | 96.62 | 97.45 |
FedProx | 86.3 | 94.23 | 95.98 | NaN | 96.63 |
FedAdam | 57.43 | 88.93 | 95.32 | NaN | 96.96 |
SCAFFOLD | 61.37 | 78.99 | 96.57 | 96.84 | 97.49 |
B.4 Performance Plots
To gauge convergence, it can be helpful to examine performance plots showing how accuracy progresses throughout federated training. Below we provide four such plots, comparing all models when training from scratch on CIFAR-10 using FedAvg and SCAFFOLD, comparing all models for the experiment in Table 4, and a Fed-ISIC run from the top performing model in Table 2, ANFR with SCAFFOLD.
Appendix C FedChest construction and additional results
C.1 Construction and hyper-parameters
@@
After pre-processing, CheXpert has twice as many samples as the other datasets, so we further split it into two clients, cxp_young and cxp_old using the median age of the patient population (63 years), leading to a total of 4 clients with train/val/test splits of (given in thousands): for CXR14, for PadChest, for cxp_old and for cxp_young. The task is multi-label classification across the 8 common classes.
@@ C.2 Additional FedChest Metrics
C.3 Batch Size Ablation Study
The absence of a performance gap between BN-ResNet and ANFR on the FedChest dataset when using FedProx (Table 2) motivates us to perform a study ablating the batch size to examine how inconsistent averaging, which is expected to happen for small batch sizes, affects results. We compare BN-ResNet and ANFR, varying the batch size while keeping all other experimental settings unchanged.
Batch Size | 16 | 32 | 64 | 128 | 256 |
BN-ResNet | 78.67+0.03 | 80.02+0.18 | 81.79+0.18 | 82.14+0.1 | 81.33+0.07 |
ANFR | 79.20+0.09 | 80.57+0.03 | 81.71+0.16 | 82.14+0.1 | 82.19+0.07 |
Table 15 shows BN-ResNet’s performance degrades more than that of ANFR for small batch sizes (16 and 32). ANFR offers significant advantages compared to BN-ResNet for small batch sizes due to the absence of BN. In the main paper hyper-parameters are tuned based on BN-ResNet’s performance; as the best BN-ResNet result is achieved with a batch size of 128, this is the one used. While using a large enough batch size can mitigate intra-client variance to a degree, we see that increasing the batch size to 256 reduces BN-ResNet’s performance, indicating diminishing returns. This reinforces that increasing batch size is ultimately not a viable solution for addressing BN’s limitations in non-IID FL, and new methods, such as ANFR, are necessary to effectively combat statistical heterogeneity.
@@ D.3 Channel attention plots of all layers
Finally, Figure 7 shows the CA weights for every CA module of SE-ResNet and ANFR.